Let.X and 7 be random losses with joint density function fr, y- for x and y non-negative integers. What is the probability that XY S 3? e-9 72 xly! A 0.016 B 0.021 co.055 D0.082 E 0.86
Answer #1

here from above f(x,y) =(e-7*7x/x!)*(e-2*2y/y!) =f(x)*f(y)

therefore x and y follows poisson distribution with parameter 7 and 2.

hence X+Y will follow poisson distribution with parameter 9.

P(X+Y<=3)=\sum_{x=0}^{3}e^{-9}9^{x}/x! =0.021 option B is correct